OpenStudy (anonymous):

An ideal gas has a density of 1.75 kg/m3 at a gauge pressure of 160 kPa. What must be the gauge pressure if a density of 1.0 kg/m3 is desired at the same temperature?

OpenStudy (anonymous):

atmospheric pressure = 101 kPa. P = absolute pressure = atm + gauge reduce density by reducing absolute P at same tmeperature density 2 / density 1 = P2/P1 = 1/1.75 P2 = (1/1.75) (101kPa + 160kPa) = 149 kPa gauge = (149-101)= 48 kPa gauge